The length of each side of an equilateral triangle is 28 centimeter. Find he altitude of the triangle. Round you answer to he nearest tenth of a centimeter.
Answer by rfer(10432)

a^2+b^2=c^2
a^2+14^2=28^2
a^2=784-196
a^2=588
a=24.3 cm